If you invest $500 today in an account that pays 6 percent interest compounded annually, how much will be in your account after two years?

1
Expert's answer
2022-03-14T19:41:52-0400

For compound interest, we use the formula below:

A=P(1+r)t

Where: A is the accumulated amount

P is the principle

R is the interest rate

T is the time

In the given question p=500, r=6% t=2 years

A=500(1+6%)2

=500(1.06)2

=500(1.1236)

=561.8

So after two years, the account will have $561.8